# Sharp Monocular View Synthesis in Less Than a Second

### v1

@@ C.1 Image Fidelity Metrics
As shown in Table 2 and Figure 4, we observe that older pointwise metrics such as PSNR and SSIM are highly sensitive to small spatial misalignments. A mere 1% translation causes PSNR to drop to 11.2 and SSIM to 0.375 – values that are surprisingly close to those obtained when comparing with a mean image (PSNR 10.7, SSIM 0.351).
@@
DISTS (↓) LPIPS (↓) PSNR (↑) SSIM (↑)
@@ C.2 Depth Estimation Uncertainty
Monocular depth estimation is fundamentally ill-posed, as multiple 3D configurations can produce the same 2D image (Poggi et al., 2020). Figure 5 illustrates this ambiguity by comparing depth predictions for an image and its mirror image – a technique similar to that used in left-right consistency for monocular depth training (Godard et al., 2017). The uncertainty map reveals that depth estimators struggle most at object boundaries and in regions with complex geometric structures, such as foliage. When these ambiguous depth estimates are used directly for view synthesis, the resulting images can exhibit visual artifacts as the network attempts to average across multiple plausible depth configurations. Our depth adjustment module, inspired by Conditional Variational Autoencoders (Sohn et al., 2015), addresses this issue by learning a scale map that refines the predicted depth during training, addressing these ambiguities in a way that optimizes for view synthesis quality rather than depth accuracy alone.
@@ D.1 Evaluation Dataset Setup
The reason for limiting the number of pairs is the slow inference speed of diffusion-based baselines. For instance, Gen3C takes 15 minutes to synthesize a new view (as a byproduct of synthesizing a video). 512 pairs already take roughly 5 days to evaluate on an A100, any larger set becomes less tractable to evaluate.
@@
ScanNet++. We sample from the nvs test split with DLSR images.
@@ D.4 Evaluation with Privileged Depth Information
For approaches where a depth proxy is available (Flash3D uses UniDepth (Piccinelli et al., 2024), ViewCrafter uses Dust3r (Wang et al., 2024b), TMPI uses DepthPro (Bochkovskii et al., 2025), Gen3C uses MoGe (Wang et al., 2024a), and SHARP uses DepthPro (Bochkovskii et al., 2025)), we align the intermediate depth representation $\mathbf{\hat{D}}$ to the ground truth $\mathbf{D}$ to derive an approximate global scale factor:
@@ D.5 Ablation Studies
Datasets. We report metrics on ScanNet++ (small scale scenes) and Tanks and Temples (large scale scenes), and display results on the real-world dataset Unsplash (Unsplash, 2022).
@@
We always incorporate color and alpha losses for appearance reconstruction. Our ablation of loss terms (Table 8 and Figure 9) show that the depth loss reduces geometry distortion, and perceptual loss brings significant improvement in inpainting quality and image sharpness; both losses result in improved metrics. While our regularizers do not move the metrics on the datasets used for ablation analysis, they qualitatively improve scenes with challenging geometry and faraway backgrounds (Figure 9). We also observe that our regularizers boost rendering speed (Table 9), which we attribute to the fact that they prevent degenerate or very large Gaussians.
@@
Table 14 evaluates the contribution of the number of Gaussians that we output from our network. We compare the full $2\times 784\times 784\approx 1.2M$ output to a $2\times$ and $4\times$ downsampled output. We see that performance of our method improves when we predict more Gaussians. This is confirmed by our qualitative results in Figure 13.
@@
$2\times 196\times 196$ 0.110 0.199 20.46 0.799 0.181 0.458 16.27 0.525
$2\times 392\times 392$ 0.077 0.160 22.00 0.822 0.140 0.425 16.23 0.525
$2\times 784\times 784$ 0.064 0.147 22.61 0.829 0.126 0.419 16.19 0.523
@@ D.6 Motion Range
In Figure 7 we study the the perceptual metrics trend against the motion values (measured by pairwise camera baseline size in meter) in our evaluation setup (see Section D.1). Experiments show that while SHARP works well, as expected, on small camera motion (<0.5<0.5 meter), it retains its quality on larger motion and performs better than most other approaches on extended motion ranges. The SOTA diffusion-based approach Gen3C only outperforms SHARP on ETH3D with motion >3>3 meter and on ScanNet++ with motion >0.5>0.5 meter. We also see that with privileged info, the quality regression over motion range can be further alleviated. In summary, quantitative analysis show that while the desired motion range is around half a meter, our approach still works reasonably well on larger camera displacement.
@@
Figure 7: Motion range analysis on the evaluation set. Shade indicates standard deviation. Unshaded data points indicate one single samples in the bin. Bins without samples are skipped, c.f. Figure 6. SHARP works consistently the best with camera baseline sizes <0.5<0.5 meter, and maintains comparable results against diffusion-based approaches on larger motion ranges. It remains the best or the second best up to 3 meters.
@@
We believe it is an interesting research direction to combine the strengths of diffusion-based approaches (synthesis of faraway content) and feed-forward models such as SHARP (interactive generation of a 3D representation that can be rendered in real time).
@@ D.7 Failure Cases
Apart from the failure of excessive motion range that exceed the operation domain, like all machine learning models, SHARP may fail under challenging scenarios. In Figure 8 we show several such examples.
@@
  • The complex reflection in water is interpreted by the network as a distant mountain, therefore the water surface is broken.
@@ D.8 Additional Qualitative Results
Figure 14: Extending the range of motion beyond nearby views, with monodepth as privileged depth information. We do not show LVSM and SVC as they cannot make use of privileged depth information, see discussions in Section D.4. ViewCrafter and Gen3C operates at a fixed aspect ratio, therefore we pad their output to match the original image resolution.

### v2

@@ C.1 Image Fidelity Metrics
As shown in Table 2 and Figure 4, we observe that older pointwise metrics such as PSNR and SSIM are highly sensitive to small spatial misalignments. A mere 1% translation causes PSNR to drop to 11.2 and SSIM to 0.375, values that are surprisingly close to those obtained when comparing with a mean image (PSNR 10.7, SSIM 0.351).
@@
DISTS↓ LPIPS↓ PSNR↑ SSIM↑
@@ C.2 Depth Estimation Uncertainty
Monocular depth estimation is fundamentally ill-posed, as multiple 3D configurations can produce the same 2D image (Poggi et al., 2020). Figure 5 illustrates this ambiguity by comparing depth predictions for an image and its mirror image, a technique similar to that used in left-right consistency for monocular depth training (Godard et al., 2017). The uncertainty map reveals that depth estimators struggle most at object boundaries and in regions with complex geometric structures, such as foliage. When these ambiguous depth estimates are used directly for view synthesis, the resulting images can exhibit visual artifacts as the network attempts to average across multiple plausible depth configurations. Our depth adjustment module, inspired by Conditional Variational Autoencoders (Sohn et al., 2015), addresses this issue by learning a scale map that refines the predicted depth during training, addressing these ambiguities in a way that optimizes for view synthesis quality rather than depth accuracy alone.
@@ D.1 Evaluation Dataset Setup
The reason for limiting the number of pairs is the slow inference speed of diffusion-based baselines. For instance, Gen3C takes 15 minutes to synthesize a new view (as a byproduct of synthesizing a video). 512 pairs already take roughly 5 days to evaluate on an A100; any larger set becomes less tractable to evaluate.
@@
ScanNet++. We sample from the nvs test split with DSLR images.
@@ D.4 Evaluation with Privileged Depth Information
For approaches where a depth proxy is available (Flash3D uses UniDepth (Piccinelli et al., 2024), ViewCrafter uses Dust3r (Wang et al., 2024), TMPI uses DepthPro (Bochkovskii et al., 2025), Gen3C uses MoGe (Wang et al., 2025), and SHARP uses DepthPro (Bochkovskii et al., 2025)), we align the intermediate depth representation $\mathbf{\hat{D}}$ to the ground truth $\mathbf{D}$ to derive an approximate global scale factor:
@@ D.5 Ablation Studies
Datasets. We report metrics on ScanNet++ (small-scale scenes) and Tanks and Temples (large-scale scenes), and display results on the real-world dataset Unsplash (Unsplash, 2022).
@@
We always incorporate color and alpha losses for appearance reconstruction. Our ablation of loss terms (Table 8 and Figure 9) shows that the depth loss reduces geometry distortion, and perceptual loss brings significant improvement in inpainting quality and image sharpness; both losses result in improved metrics. While our regularizers do not move the metrics on the datasets used for ablation analysis, they qualitatively improve scenes with challenging geometry and faraway backgrounds (Figure 9). We also observe that our regularizers boost rendering speed (Table 9), which we attribute to the fact that they prevent degenerate or very large Gaussians.
@@
Table 14 evaluates the contribution of the number of Gaussians that we output from our network. We compare the full $2\times 768\times 768\approx 1.2M$ output to a $2\times$ and $4\times$ downsampled output. We see that performance of our method improves when we predict more Gaussians. This is confirmed by our qualitative results in Figure 13.
@@
$2\times 192\times 192$ 0.110 0.199 20.46 0.799 0.181 0.458 16.27 0.525
$2\times 384\times 384$ 0.077 0.160 22.00 0.822 0.140 0.425 16.23 0.525
$2\times 768\times 768$ 0.064 0.147 22.61 0.829 0.126 0.419 16.19 0.523
@@ D.6 Motion Range
In Figure 7 we study the perceptual metrics trend against the motion values (measured by pairwise camera baseline size in meters) in our evaluation setup (see Section D.1). Experiments show that while SHARP works well, as expected, on small camera motion (<0.5<0.5 meters), it retains its quality on larger motion and performs better than most other approaches on extended motion ranges. The SOTA diffusion-based approach Gen3C only outperforms SHARP on ETH3D with motion >3>3 meters and on ScanNet++ with motion >0.5>0.5 meters. We also see that with privileged info, the quality regression over motion range can be further alleviated. In summary, quantitative analysis shows that while the desired motion range is around half a meter, our approach still works reasonably well on larger camera displacement.
@@
Figure 7: Motion range analysis on the evaluation set. Shade indicates standard deviation. Unshaded data points indicate a single sample in the bin. Bins without samples are skipped, cf. Figure 6. SHARP works consistently the best with camera baseline sizes <0.5<0.5 meters, and maintains comparable results against diffusion-based approaches on larger motion ranges. It remains the best or the second best up to 3 meters.
@@
We believe it is an interesting research direction to combine the strengths of diffusion-based approaches (synthesis of faraway content) and feedforward models such as SHARP (interactive generation of a 3D representation that can be rendered in real time).
@@ D.7 Failure Cases
Apart from the failure of excessive motion ranges that exceed the operation domain, like all machine learning models, SHARP may fail under challenging scenarios. In Figure 8 we show several such examples.
@@
  • The complex reflection in water is interpreted by the network as a distant mountain; therefore, the water surface appears broken.
@@ D.8 Additional Qualitative Results
Figure 14: Extending the range of motion beyond nearby views, with monodepth as privileged depth information. We do not show LVSM and SVC as they cannot make use of privileged depth information, see discussions in Section D.4. ViewCrafter and Gen3C operate at a fixed aspect ratio, therefore we pad their output to match the original image resolution.